Find x and y:
x+y=7 and x*y=11

Answer by stanbon(75887)

Find x and y:
x+y=7 and x*y=11
--------------------
Substitute for "y" to get:
x(7-x) = 11
-x^2+7x = 11
---
-x^2 + 7x - 11 = 0
---
x^2 - 7x + 11 = 0
---
x = [7 +- sqrt(49 -4*11)]/2
---
x = [7 +- sqrt(5)]/2
x = [7+sqrt(5)]/2 or x = [7-sqrt(5)]/2
x = (7/2) + (1/2)sqrt(5) or x = (7/2) - (1/2)sqrt(5)
-----------------
Since y = 7-x
y = (7/2)-(1/2)(sqrt)5 or y = (7/2) + (1/2)sqrt(5)
=====================================================
Cheers,
Stan H.
y = 7 -